OpenStudy (ssprague_77):

4y - 6y + 9y = -2 What is the solution set of this equation? SOLVE!

OpenStudy (trojanpoem):

(4-6+9)y = -2 7y = -2 y = -2/7

OpenStudy (ssprague_77):

THANK YOU!! Your a life saver! I am in a test

563blackghost (563blackghost):

Please do not post test question. We can not help with test, because a test...tests your knowledge of what you learned in the unit.

OpenStudy (ssprague_77):

A = 3(B + C) solve for B....

OpenStudy (trojanpoem):

@563blackghost , SHET , I helped him cheat in an exam xD

563blackghost (563blackghost):

Well you showed the steps and you didn't know so your totally fine :)